OpenStudy (anonymous):

A football player rushed 120 yards. Which measurment represents the same measurments?

OpenStudy (magbak):

IN what feet or inches.

OpenStudy (anonymous):

its on feet

OpenStudy (magbak):

It is 120 time 3 and that is 360 feet.

OpenStudy (anonymous):

thanks

OpenStudy (magbak):

You are welcome

OpenStudy (whpalmer4):

120 yards in feet: 1 yard = 3 feet 120 yards * (3 feet / 1 yard) = 360 feet notice that the yards cancel out. If you write the conversion incorrectly, the units do not cancel correctly and you have a clue that you've multiplied where you should have divided, or not applied enough conversion factors, etc.

OpenStudy (anonymous):

Rashelle bought a square picture frame to hang on her wall that covers one hundred twenty-one inches squared of wall area. What is the side length of the picture frame?

OpenStudy (anonymous):

help?

OpenStudy (magbak):

It is 11 inches

OpenStudy (whpalmer4):

Area of a square or rectangle is given by \[A = l*w\] You have A = 121 sq inches, and length = width \[121 \text{ in}^2 = w^2\]Take square of both sides \[\sqrt{121 \text{ in}^2} = w = 11 \text{ in}\]

OpenStudy (whpalmer4):

make that "take square root"